# The Hands-Up Problem and How to Deal With It: Secondary School Teachers’ Experiences of Debugging in the Classroom

### v1

@@ 6 Conclusions and Further Work
We have reported secondary school teachers’ experiences of debugging in the classroom, with a specific focus on text-based programming. General experiences and perspectives of debugging, as well as teachers’ observations of students debugging, were all discussed and analysed through thematic analysis.
@@
The themes we identified notably related to the common reliance of students on the teacher for debugging support, corroborating with some previous work (Bagge, 2015; Sentance and Csizmadia, 2017; Michaeli and Romeike, 2019a; Uppal, 2020; Kerslake, 2024). Due to this reliance being reflected through lots of raised hands in the classroom, we conceptualised our themes through the phenomenon of the ‘hands-up problem’. We define this as the in-class difficulty deriving from the common student tendency to raise their hand when struggling to complete a task. There are a number of drivers for the frequent hand-raising that contribute to a busy classroom atmosphere. This has generally negative consequences; students are often left stuck and frustrated, and teachers flustered. Some teachers, however, had developed counters to the hands-up problem, usually through experience. These consisted of general classroom management techniques as well as debugging-specific teaching approaches. Over time, these help students to become less reliant on the teacher and more resilient to errors in their programs.
@@
Based on these results, we suggest three areas of further work. First, developing a conceptualisation of debugging PCK would be a useful way of synthesising research on teacher-centric perspectives of debugging in the classroom. This should include resources and approaches developed by both researchers and teachers, which would allow for iterative testing and refinement in a classroom context. Teachers’ debugging PCK could then be improved through professional development. Further development and expansion of debugging-related PCK is a second area of further work we recommend.
Finally, understanding what has helped confident computing teachers to foster teacher-independent debugging would be useful. Whether teachers attribute this to experience, professional development, or intentional reflection on their teaching practice would play an important role in the way that debugging PD is structured and engaged with. This development of counters to the hands-up problem, illustrated by the far-left box in Figure 1, is not something our thematic analysis identified. Further capturing the experiences of teachers will help to improve their, and their students’, experiences with debugging.

### v3

@@ 6 Conclusions and Further Work
This paper has reported secondary school teachers’ experiences with the teaching and learning of debugging in their programming lessons, with a specific focus on text-based programming. Through thematically analysed semi-structured interviews, we report teachers’ strategies for supporting students with debugging, their observations of how students engage with the debugging process, and some challenges they face related to debugging.
@@
A particular trend among our themes was the common student reliance on the teacher when encountering and struggling to resolve errors. This corroborates with previous reports of computing teachers’ classroom experiences (CombattingLearnedHelplessness; ComputingInCurriculumChallengesStrategies; CurrentPerspectivesDebugging; LackPerserveranceAutonomyDebugging; StumpTheTeacher). We conceptualised our results through the hands-up problem, the in-class difficulty deriving from the common student tendency to raise their hand when struggling to complete a task. There are several drivers for the frequent hand-raising that contributes to the hands-up problem, which have generally negative consequences. Students are often left stuck and frustrated, and teachers flustered. Usually through experience, some teachers developed counters to the hands-up problem. These consisted of general classroom management techniques and debugging-specific teaching approaches. Over time, these counters help students to become less reliant on the teacher and more resilient to errors in their programs.
This paper presents the hands-up problem as an initial conceptualisation of a significant challenge teachers face in programming lessons as a result of student struggle with debugging. This conceptualisation should be developed through further studies of teachers’ experiences with teaching programming in the classroom. Where previous studies have used interviews (CurrentPerspectivesDebugging) and simulations of teacher-student interactions (DoINeedToKnowIfTheTeacher; SupportingStudentsScienceInquiry; TeacherDiagnosticInterventionSkills), methods such as classroom observations would help to paint a more well-rounded picture of teachers’ lived experiences. This work is increasingly important to conduct given the increasing number of school curricula mandating the teaching of text-based programming (e.g., (UKSecondaryComputingCurriculum; BritishColumbiaADSTCurriculum; PolandCSCurriculumChallengesSolutions; ProgrammingSkillsInLatinAmerica; Swedish2024Curriculum)).
@@
Developing a model for debugging PCK is another area of further work we recommend. This would be a useful way of synthesising research into beginner debugging education at a school level and beyond. A debugging PCK should include approaches developed by both researchers and teachers, which could be iteratively tested and refined in the classroom. This could then inform further work on debugging-specific professional development and pedagogy.
Finally, further investigation into pedagogical approaches that facilitate positive experiences with learning to debug should be conducted. This is already happening, but future research should investigate these approaches on a larger scale over longer periods of time to better understand their effectiveness (DebuggingInterventionLitReview). Supporting teachers’ debugging pedagogies will help to cultivate classroom environments where errors are confidently handled rather than feared.
{acknowledgement}
[title=Acknowledgments] We are very grateful to all of the teachers we interviewed for sharing their invaluable experiences. Without their participation, this research would not have been possible.
The first author gives all praise to the Lord Jesus Christ for blessing him with the wisdom and opportunity to conduct this research.